please, I need clear writing if you choose to do it by hand.

thanks,

2. A device consists of two redundant subunits; only one needs to be working for the device to function. The first subunit has two parts Wi and W2, and the second has two parts W3 and W4. In each subunit, both parts must function for the subunit to work. The four parts are independent identical and equally reliable, with probability p- 0.9 of functioning. (a) What is the overall probability the device functions correctly? (b) Suppose you have available two ultra-reliable parts, each with p-0.99 of working. Which two parts in the device should you replace with these in order to maximize device reliability? Justify your answer by calculating the resulting reliability and explain your choice,
Answer #1

(a)

W1 & W2 are in parallel with each p = 0.9

Reliability of W1 W2 = 1 - (1 - 0.9)2 = 0.99

W3 & W4 are in parallel with each p = 0.9

Reliability of W3 W4 = 1 - (1 - 0.9)2 = 0.99

W1W2 and W3W4 are in series.

So,

Overall probability the device functions correctly = 0.99 X 0.99 = 0.9801

(b)

Case 1:

W1 = 0.99, W2 =0.99, W3 =0.9, W4 = 0.9

Reliability of W1 W2 = 1 - (1 - 0.99)2 = 0.9999

Reliability of W3 W4 = 1 - (1 - 0.9)2 = 0.99

So,

Overall probability the device functions correctly = 0.9999 X 0.99 = 0.9899

Case 2:

W1 = 0.9, W2 =0.9, W3 =0.99, W4 = 0.99

Reliability of W1 W2 = 1 - (1 - 0.9)2 = 0.99

Reliability of W3 W4 = 1 - (1 - 0.99)2 = 0.9999

So,

Overall probability the device functions correctly = 0.99 X 0.9999 = 0.9899

Case 3:

W1 = 0.99, W2 =0.9, W3 =0.99, W4 = 0.9

Reliability of W1 W2 = 1 - (1 - 0.99)(1-0.9) = 0.999

Reliability of W3 W4 = 1 - (1 - 0.99) (1-0.9)= 0.999

So,

Overall probability the device functions correctly = 0.999 X 0.999= 0.9980

Case 4:

W1 = 0.9, W2 =0.99, W3 =0.9, W4 = 0.99

Reliability of W1 W2 = 1 - (1 - 0.9)(1-0.99) = 0.999

Reliability of W3 W4 = 1 - (1 - 0.9) (1-0.99)= 0.999

So,

Overall probability the device functions correctly = 0.999 X 0.999= 0.9980

We note from the above:

Cases 3 & 4 each give overall reliability = 0.9980 > 0.9899 for Cases 1 & 2.

So,

we should replace parts 1 & 3 each with p = 0.99 in order to maximize device reliability

OR

we should replace parts 2 & 4 each with p = 0.99 in order to maximize device reliability